OpenStudy (anonymous):

The probability that Shelly will go to a movie (event A) on Friday is 0.78, and the probability that Danielle will go to a movie (event B) the same day is 0.63. The probability that Shelly will go to a movie, given that Danielle goes to a movie, is 0.87. Which statement is true?

OpenStudy (anonymous):

Events A and B are independent because P(A|B) P(A). Events A and B are independent because P(A|B) = P(B). Events A and B are dependent because P(A|B) = P(A) P(B). Events A and B are dependent because P(A|B) P(A). Events A and B are dependent because P(A|B) = P(A).